# Accurate Network Traffic Matrix Prediction via LEAD: an LLM-Enhanced Adapter-Based Conditional Diffusion Model

### v1

@@ 4.4 Main Results and Analysis
(a) RMSE Comparison with STGNNs on Abilene dataset.
(b) RMSE Comparison with STGNNs on GÉANT dataset.
(c) RMSE Comparison with Multi-Scale Methods on Abilene dataset.
(d) RMSE Comparison with Multi-Scale Methods on GÉANT dataset.
@@ 4.5 Ablation Studies
TABLE III: Ablation studies of different components.
@@
5 Conclusion
In this paper, we propose LEAD, a framework that synergizes the semantic reasoning capability of a frozen LLM (via parameter-efficient adapters) with the probabilistic modeling power of conditional diffusion for TM forecasting. By adopting a generative perspective rather than purely regressive prediction, LEAD mitigates the over-smoothing issue of conventional baselines and reduces dependence on static graph topologies. Experiments on Abilene and GÉANT validate that semantic-guided diffusion achieves higher-fidelity forecasts under bursty and non-stationary dynamics. These results support foundation-model-empowered predictive networking as a practical building block for AI-native edge operation, especially when efficient adaptation and deployment under resource constraints are required.
Despite these promising results, two limitations remain to be addressed. First, the iterative denoising mechanism inherently limits inference efficiency compared to one-step baselines. Second, the reliance on a frozen, lightweight LLM, select to maintain computational balance, constrains deep domain adaptation, leaving the full potential of LLM-driven reasoning in network physics largely underexplored.
@@ 5 Conclusion
(2) unlocking the full potential of the semantic core by exploring parameter-efficient fine-tuning like LoRA [39] on larger-scale models to better capture intricate network dynamics. We also aim to extend this framework to broader tasks, such as network anomaly detection and root cause analysis.

### v2

@@ IV-D Main Results and Analysis
(a) RMSE Comparison of LEAD with STGNNs on the Abilene dataset.
(b) RMSE Comparison of LEAD with STGNNs on the GÉANT dataset.
(c) RMSE Comparison of LEAD with Multi-Scale methods on the Abilene dataset.
(d) RMSE Comparison of LEAD with Multi-Scale methods on the GÉANT dataset.
@@ IV-E Ablation Studies
TABLE III: Ablation studies of different components, showing RMSE and MAE values on the Abilene and GÉANT datasets.
@@
V Conclusion
In this paper, we propose LEAD, a competitive model that synergizes the semantic reasoning capability of a frozen LLM (via parameter-efficient adapters) with the probabilistic modeling power of conditional diffusion for TM forecasting. By adopting a generative perspective rather than purely regressive prediction, LEAD mitigates the over-smoothing issue of conventional baselines and reduces dependence on static graph topologies. Experiments on Abilene and GÉANT validate that semantic-guided diffusion achieves higher-fidelity forecasts under bursty and non-stationary dynamics. These results support foundation-model-empowered predictive networking as a practical building block for AI-native edge operation, especially when efficient adaptation and deployment under resource constraints are required.
Despite these promising results, two limitations remain to be addressed. First, the iterative denoising mechanism inherently limits inference efficiency compared to one-step baselines. Second, the reliance on a frozen, lightweight LLM, selected to maintain computational balance, constrains deep domain adaptation, leaving the full potential of LLM-driven reasoning in network physics largely underexplored.
@@ V Conclusion
(2) unlocking the full potential of the semantic core by exploring parameter-efficient fine-tuning like Low-Rank Adaptation (LoRA) [42] on larger-scale models to better capture intricate network dynamics. We also aim to extend this framework to broader tasks, such as network anomaly detection and root cause analysis.